Diane is 23 year older than her daughter Amy in 6 year Diane will be twice as old as Amy how old are they now
Answer by ankor@dixie-net.com(15650)

Let d = Diane's present age
Let a = Amy's present age
:
Write an equation for each statement:
:
Diane is 23 year older than her daughter, Amy.
d = a + 23
;
In 6 years Diane will be twice as old as Amy.
d + 6 = 2(a+6)
d + 6 = 2a + 12
d = 2a + 12 - 6
d = 2a + 6
Replace d with (a+23) from the 1st statement
a + 23 = 2a + 6
23 - 6 = 2a - a
17 = a; Amy is 17
:
I'll let you find Diane's age